OpenStudy (cookiemonster18):

Graph the direct variation equation. y=-1/2x

OpenStudy (danjs):

to graph a line, you need 2 points

OpenStudy (danjs):

plot 2 points, draw the line through them

OpenStudy (cookiemonster18):

A (the first graph)? @DanJS

OpenStudy (danjs):

hi, the thing for sure goes through the origin, if x=0, y is also 0

OpenStudy (danjs):

y = (-1/2)*x is just the line equation , y = m*x + b with b=0 and slope m=-1/2

OpenStudy (danjs):

the first graph has a positive slope

OpenStudy (danjs):

second graph has slope = rise/run = -2/1=-2 not -1/2

OpenStudy (danjs):

the last graph is y=-1/2 *x every 2 increase oin x results in -1 decrease in y,

OpenStudy (cookiemonster18):

im confused now sorry. but D (the last graph) is -1/2??